OpenStudy (anonymous):

another...lim x->0, x+xcosx/sinxcosx?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[\frac{x(1+\cos(x))}{\sin(x)\cos(x)}\] the break in two parts as before

OpenStudy (anonymous):

i got that far...let me c if the next step jumps out at me

OpenStudy (anonymous):

think about a limit that you know (hint, sine) and a limit that is obvious

OpenStudy (anonymous):

haha its not

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

limit is 0, but its not clicking...

OpenStudy (anonymous):

that is because it is not zero

OpenStudy (anonymous):

oh man that i am really lost now..i thought the lim of sine at 0 is 0?

OpenStudy (anonymous):

\[\frac{x}{\sin(x)}\times \frac{1+\cos(x)}{cos(x)}\]

OpenStudy (anonymous):

x/sinx=1 righ?

OpenStudy (zarkon):

as x goes to zero...yes

OpenStudy (anonymous):

still lost on how the 2 comes out as the answer?

OpenStudy (anonymous):

does 1+cosx/cosx=2?

OpenStudy (zarkon):

cos(0)=1

OpenStudy (zarkon):

(1+1)/1=2

OpenStudy (anonymous):

1+cos(x) is 1+1?

OpenStudy (anonymous):

got it!!!!!!!

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{1+\cos(x)}{\cos(x)}=\frac{1+1}{1}=2\]

OpenStudy (anonymous):

thanks zarkon and satel!!!